# Focused Transformer: Contrastive Training for Context Scaling

### v1

@@ Abstract
Large language models have an exceptional capability to incorporate new information in a contextual manner. However, the full potential of such an approach is often restrained due to a limitation in the effective context length. One solution to this issue is to endow an attention layer with access to an external memory, which comprises of (key, value) pairs. Yet, as the number of documents increases, the proportion of relevant keys to irrelevant ones decreases, leading the model to focus more on the irrelevant keys. We identify a significant challenge, dubbed the distraction issue, where keys linked to different semantic values might overlap, making them hard to distinguish. To tackle this problem, we introduce the Focused Transformer (FoT), a technique that employs a training process inspired by contrastive learning. This novel approach enhances the structure of the (key, value) space, enabling an extension of the context length. Our method allows for fine-tuning pre-existing, large-scale models to lengthen their effective context. This is demonstrated by our fine-tuning of $3B$ and $7B$ OpenLLaMA checkpoints. The resulting models, which we name LongLLaMA [^1] We release the checkpoints and the source code of LongLLaMA [Uncaptioned image], see also our colab., exhibit advancements in tasks requiring a long context. We further illustrate that our LongLLaMA models adeptly manage a $256k$ context length for passkey retrieval.
@@ 1 Introduction
Language models have served as a catalyst for substantial advancements in several areas, including natural language processing (Radford et al., 2019; Brown et al., 2020), code generation (Chen et al., 2021; Li et al., 2022), quantitative reasoning (Lewkowycz et al., 2022) and theorem proving (Polu and Sutskever, 2020; Jiang et al., 2022; Mikuła et al., 2023). One of the central challenges with language models is the effective incorporation of extensive new knowledge. The common practice of fine-tuning the model is not only resource-intensive and complex to manage, but it also does not always clearly indicate how to incorporate new knowledge. For example, fine-tuning on a text such as “Alice in Wonderland” does not equip the model to answer questions about the story itself, but rather it trains the model to predict the next token or complete masked sentences. A promising alternative – integrating the new knowledge within the context – doesn’t require training but is considerably restricted by the model’s effective context length. For this method to work with large knowledge databases, the model needs to manage a context length extending to millions of tokens.
@@
We propose the Focused Transformer (FoT), an innovative technique developed explicitly to address this issue. The Focused Transformer permits a subset of attention layers to access an external memory of (key, value) pairs through the k-nearest neighbors (kNN) algorithm, akin to the method used in (Wu et al., 2022). This mechanism effectively extends the total context length. The distinctive aspect of the Focused Transformer is its training procedure, drawing from contrastive learning. This method addresses the distraction issue and facilitates larger memory capacities. Specifically, during the training phase, we deliberately expose the memory attention layers to both relevant and irrelevant keys (like negative samples from unrelated documents). This strategy incentives the model to differentiate keys connected with semantically diverse values, thereby enhancing their structure.
@@
2. We develop the Focused Transformer (FoT), designed to alleviate the distraction issue. FoT includes a unique training objective that improves the (key, value) structure, enabling the use of extensive external memory and k-nearest neighbors lookup to scale the context length.
3. Our method is simple to implement, and it provides the benefit of augmenting existing models with memory without modifying their architecture, facilitated by cost-effective fine-tuning. We demonstrate this on the $3B$ and $7B$ OpenLLaMA checkpoints. The resulting models, named LongLLaMAs, display enhancements on tasks that benefit from increasing the number of few-shot demonstrations in the extended context, such as TREC (Li and Roth, 2002; Hovy et al., 2001) and WebQS (Berant et al., 2013). We also prove that for passkey retrieval Mohtashami and Jaggi (2023), our LongLLaMA models successfully handle a $256k$ context length.
@@ 2 Related work
A multitude of approaches have been developed to increase the context length of transformers, mostly focusing on alleviating the quadratic complexity of the attention computation. For instance, Transformer-XL (Dai et al., 2019) caches the previous context and enables the linear extension of context with the number of layers. Longformer (Beltagy et al., 2020) employs an attention mechanism that allows tokens to attend to distant tokens sparsely, reducing the computational complexity. BigBird (Zaheer et al., 2020), LongT5 (Guo et al., 2021), and (Dao et al., 2022) also use sparse attention to handle long sequences. Hourglass (Nawrot et al., 2021) downsamples activations in intermediate layers to reduce computation and enable longer contexts. COLT5 (Ainslie et al., 2023) proposes conditional computation to save memory and enable larger contexts. Memorizing Transformer (Wu et al., 2022) uses kNN lookup to pick up the most relevant tokens, which might also be seen as a way to reduce the computational complexity of attention. Our work adheres to this approach and aims to train a key space that handles longer attention context length (e.g., by mitigating the distraction issue) and, thus, has better long-context capabilities.
Fine-tuning LLMs for longer context
@@
Prior works such as RETRO (Borgeaud et al., 2022) (RETROfitting) and Memorizing Transformer (Wu et al., 2022) have demonstrated a promising path for fine-tuning existing LMs to add new capabilities without the need to retrain the entire model. More recently, a number of works have explored fine-tuning LLaMA to extend its context length. Landmark attention (Mohtashami and Jaggi, 2023) proposes a compression scheme of LLM’s context into landmarks, increasing the context length of LLaMA-7B to $32K$. Position Interpolation (PI, (Chen et al., 2023) and (kaiokendev, 2023)) introduces a modification to the rotary positional encoding scheme that enables fine-tuning for $32K$ context. In contrast to this work, our method does not rely on positional encodings, following the findings from (Haviv et al., 2022). Removing positional encoding in memory allows us to extrapolate to $256k$ tokens, although the model was only trained on sequences up to $8K$, yielding theoretically unbounded context length.
@@
Contrastive learning aims to learn good representations by comparing positive and negative examples. CLIP (Radford et al., 2021) and SimCLR (Chen et al., 2020) are two popular contrastive learning methods that have achieved state-of-the-art performance in the image domain. During contrastive pre-training, negative examples are kept in the same batch to learn to distinguish them from positive examples. Scaling the batch size in contrastive learning has been demonstrated to enhance the quality of representations, as shown in (Gao et al., 2021b). It has been suggested (Gao et al., 2019) that the embedding space in language modeling suffers from degeneracy, where embeddings are tightly packed in a narrow cone, making it difficult to distinguish between them. TRIME (Zhong et al., 2022) proposes a training approach designed for training LMs with memory augmentation, which uses in-batch negatives to improve the quality of representations. The main difference between this and our approach is that we incorporate negatives into the memory attention layer instead of interpolating in the output layer.
@@ 3 FoT: Focused Transformer
Our method, the Focused Transformer (FoT), is a simple plug-and-play extension of transformer models and can be used both to train new models or fine-tune existing, possibly large, models with longer context. To this end, FoT uses memory attention layers and the crossbatch training procedure. Memory attention layers enable the model to retrieve information from the external memory at inference time, effectively extending the context. The crossbatch training procedure biases the model to learn $(key,value)$ representations, which are easy to use by a memory attention layer. See Figure 2 for an overview of the FoT architecture and Appendix F for pseudocode.
Figure 2: The Focused Transformer overview. During inference, a memory attention layer (green) uses external memory of $(key,value)$ pairs via kNN lookup, which effectively extends its context length. This layer is trained using crossbatch. Namely, the tokens from the current context $C_{curr}$ attend in a differentiable way (Att + $\nabla$) to the previous context $C_{prev}$ of the same document and, importantly, $d-1$ contexts of other documents. The latter serve as ’negative’ examples intended to better shape the $(key,value)$ space.
@@ 3.1 Memory attention layers
Memory attention layers $\mathcal{L}$ are endowed with access to an external memory database during inference. Namely, each query in $\ell\in\mathcal{L}$ attends to preceding keys from the local context and the top $k$ most matching keys from memory. The memory keys are ranked by the inner product with the query and retrieved using the kNN search algorithm. We use the exact kNN search implemented in FAISS (Johnson et al., 2017). The memory is populated incrementally with $(key,value)$ pairs processed by $\ell$ beforehand. Our memory attention layer design is closely related to (Wu et al., 2022), we follow most of its design choices, except for the gating, which we replace with a simpler mechanism, which turns out to be more effective in our applications. See details in Section 5.6.3 and Appendix B.2. We remove positional encodings in memory layers in all our models except LongLLaMAs. This allows LongLLaMA checkpoints to be a drop-in replacement for LLaMA checkpoints.
@@ 3.2 Crossbatch training procedure
We do not use external memory during training. This has two important consequences. One, the operation is fully differentiable, and thus, we improve all the $(key,value)$ pairs in $p^{\delta}$. Two, the procedure is easy to implement; it does not require any additional loss (i.e., uses the standard transformer training objective) and is done on the level of the data loading pipeline and a minor self-attention change. The only new hyperparameter is $d$, which prescribes the ratio of positive to negative samples. Typically, we find it beneficial to start with small $d\leq 8$ (otherwise, the model tends to ignore the previous local context) and later switch to bigger values, say $d\geq 64$. Appendix B.3 provides more details about the method.
@@ 3.3 The distraction issue
In this section, we conceptualize what we call the distraction issue and hypothesize it is one of the key problems in using large memory databases. Namely, during the standard training, the model is not incentivized to distinguish the keys from different documents. We measure that the attention mass is evenly spread on the related and unrelated documents; see Figure 3. More precisely, for a document $\delta$, let $w_{ij}$ be the softmax weights related to $p^{\delta}_{ij}$ constructed as described in Section 3.2. We define the positive attention mass as $r_{d}:=\sum_{j}w_{1j}/\sum_{i=1}^{d}\sum_{j}w_{ij}$. We observe that $r_{d}\approx 1/d$, which can be interpreted as the fact that the attention is equally distracted by the positive (coming from the current document at $i=1$) and negative keys. This is an undesirable property since when scaling the memory, the attention becomes increasingly distracted. We show that the crossbatch mostly alleviates the distraction issue, resulting in a focused attention. More information can be found in Appendix B.4. In Section 5.4, we also show that the distraction issue has a harmful effect on metrics like perplexity.
@@ 4.2 Context length extrapolation on the passkey retrieval task
There is an important info hidden inside a lot of irrelevant text. Find it and memorize them. I will quiz you about the important information there.
<prefix filler by continuously repeating: The grass is green. The sky is blue. The sun is yellow. Here we go. There and back again.>
The pass key is <PASS KEY>. Remember it. <PASS KEY> is the pass key.
<suffix filler>
What is the pass key? The pass key is
Prompt Format used in the passkey retrieval task, copied from (Mohtashami and Jaggi, 2023).
@@ 4.5 Performance on short-context tasks
Fine-tuning for longer contexts could hurt performance on the original context length ($2K$), as the training data distribution changes. We show that this is not the case for the LongLLaMA models by evaluating them using the LM Evaluation Harness library (Gao et al., 2021a). On most tasks, the performance is kept intact; see Appendix A.4 for details, and Table 3 for the average scores. This also confirms that LongLLaMAs could be used as a drop-in replacement of LLaMA models as they are compatible with the original LLaMA inference code.
Table 3: Comparsion with OpenLLaMA models on the original context length of $2K$. We provide the average score calculated on Language Model Evaluation Harness (Gao et al., 2021a). Detailed results can be found in Table 7 in Appendix A.4.
lm-eval OpenLLaMA 3B LongLLaMA 3B OpenLLaMA 7B LongLLaMA 7B
Avg score 0.53 0.53 0.55 0.55
@@ 5.1 Experimental setup
Architecture For experiments described in this section we use decoder-only Transformer (Vaswani et al., 2017) models with $12$ layers and $184M$ parameters (unless stated otherwise, e.g., in fine-tuning experiments in Section 5.3 we scale to $1.2$B). Following Wu et al. (2022); we pick $\ell=8$ as the memory attention layer. We tune $k=128$, the number of top keys retrieved by kNN. In most experiments, we start training with a small crossbatch dimension $d\leq 8$ and switch to $d\geq 64$ after some training. For more details about the architecture and hyperparameters, see Appendix B and Appendix C.
Evaluation We distinguish two evaluation settings: single-document (abbreviated to single-doc) and multi-document (abbreviated to multi-doc). The single-doc setting is typically used for evaluating models that process long contexts. Here, we clear the memory for each new document, ensuring that only the current document is available in the context. The multi-doc setting retains memory across multiple documents without resets. This scenario is akin to a long-context retrieval task where the model must focus on tokens from useful documents without getting distracted by irrelevant ones.
Datasets We evaluate on the following long-context language modeling datasets: PG-19 (English books), arXiv (mathematical papers), GitHub (code), and Isabelle (formal proofs). PG-19 (Rae et al., 2019) is a large dataset of English-language books published prior to 1919, sourced from the Project Gutenberg archive. This dataset is a well-established benchmark for evaluating long-context language models (Sun et al., 2021). The arXiv dataset contains source of papers labeled as "Mathematics" that were obtained by downloading articles through the arXiv Bulk Data Access. The token count per paper in this dataset is comparable to that of a book in PG19. For details on the remaining datasets, refer to Appendix G.
5.2 Scaling of the context length to $16$M
We use a synthetic dictionary lookup task to check whether the model trained with our method can utilize a large memory to extend its context length. In this task, the model is first provided with $k_{i}:v_{i}$ mappings and then asked what value is associated with a particular key. We train models using documents of length $512$. The first half of each document defines keys and values associated with them, whereas the second consists of questions about values associated with keys defined in the first half. Details about the task can be found in Appendix D. For this task, we use smaller $37M$ parameter models.
For FoT, we use a local context of $256$, thus the model needs to use the memory attention layer to answer the questions correctly. We start with $d=1$ and increase to $d=128$ as soon as the model is able to reach $98\%$ training accuracy. During the inference, we use $k=32$ (the number of keys retrieved by kNN). Figure 4 shows that FoT, after $5$k steps of training, can effectively utilize memory consisting of $16$M tokens achieving accuracy above $92\%$.
As a baseline, we use a standard transformer model trained with the context length of $512$. In evaluation, we test different local context lengths, which quickly leads to very poor results.
[…]

### v2

@@ Abstract
Large language models have an exceptional capability to incorporate new information in a contextual manner. However, the full potential of such an approach is often restrained due to a limitation in the effective context length. One solution to this issue is to endow an attention layer with access to an additional context, which comprises of (key, value) pairs. Yet, as the number of documents increases, the proportion of relevant keys to irrelevant ones decreases, leading the model to focus more on the irrelevant keys. We identify a significant challenge, dubbed the distraction issue, where keys linked to different semantic values might overlap, making them hard to distinguish. To tackle this problem, we introduce the Focused Transformer (FoT), a technique that employs a training process inspired by contrastive learning. This novel approach enhances the structure of the (key, value) space, enabling an extension of the context length. Our method allows for fine-tuning pre-existing, large-scale models to lengthen their effective context. This is demonstrated by our fine-tuning of $3B$ and $7B$ OpenLLaMA checkpoints. The resulting models, which we name LongLLaMA [^1] We release the checkpoints and source code of LongLLaMA [Uncaptioned image], see also our colabs., exhibit advancements in tasks requiring a long context. We further illustrate that our LongLLaMA models adeptly manage a $256k$ context length for passkey retrieval.
@@ 1 Introduction
Language models have served as a catalyst for substantial advancements in several areas, including natural language processing (Radford et al., 2019; Brown et al., 2020), code generation (Chen et al., 2021; Li et al., 2022), quantitative reasoning (Lewkowycz et al., 2022) and theorem proving (Polu and Sutskever, 2020; Jiang et al., 2022; Mikuła et al., 2023). One of the central challenges with language models is the effective incorporation of extensive new knowledge. The common practice of fine-tuning the model is not only resource-intensive and complex to manage, but it also does not always clearly indicate how to incorporate new knowledge. For example, fine-tuning on a text such as “Alice in Wonderland” does not equip the model to answer questions about the story itself, but rather it trains the model to predict the next token or complete masked sentences. A promising alternative – integrating the new knowledge within the context – doesn’t require training but is considerably restricted by the model’s effective context length. For this method to work with large knowledge databases (like large code repositories), the model needs to manage a context length extending to millions of tokens.
@@
We propose the Focused Transformer (FoT), an innovative technique developed explicitly to address this issue. The Focused Transformer permits a subset of attention layers to access an additional context of (key, value) pairs through the k-nearest neighbors (kNN) algorithm, akin to the method used in (Wu et al., 2022). This mechanism effectively extends the total context length. The distinctive aspect of the Focused Transformer is its training procedure, drawing from contrastive learning. This method addresses the distraction issue and facilitates larger context capacities. Specifically, during the training phase, we deliberately expose the chosen subset of attention layers to both relevant and irrelevant keys (like negative samples from unrelated documents). This strategy incentives the model to differentiate keys connected with semantically diverse values, thereby enhancing their structure.
@@
2. We develop the Focused Transformer (FoT), designed to alleviate the distraction issue. FoT includes a unique training objective that improves the (key, value) structure, enabling the use of extensive additional context and k-nearest neighbors lookup to scale the context length.
3. Our method is simple to implement, and it provides the benefit of extending model context without modifying the architecture, facilitated by cost-effective fine-tuning. We demonstrate this on the $3B$ and $7B$ OpenLLaMA checkpoints. The resulting models, named LongLLaMAs, display enhancements on tasks that benefit from increasing the number of few-shot demonstrations in the extended context, such as TREC (Li and Roth, 2002; Hovy et al., 2001) and WebQS (Berant et al., 2013). We also prove that for passkey retrieval Mohtashami and Jaggi (2023), our LongLLaMA models successfully handle a $256k$ context length.
@@ 2 Related work
A multitude of approaches have been developed to increase the context length of transformers, mostly focusing on alleviating the quadratic complexity of the attention computation. For instance, Transformer-XL (Dai et al., 2019) caches the previous context and enables the linear extension of context with the number of layers. Longformer (Beltagy et al., 2020) employs an attention mechanism that allows tokens to attend to distant tokens sparsely, reducing the computational complexity. BigBird (Zaheer et al., 2020), LongT5 (Guo et al., 2021), and (Dao et al., 2022) also use sparse attention to handle long sequences. Different efficiency considerations have been studied in (Kaddour et al., 2023), showing that they lead to limited gains. Hierarchical transformers (Nawrot et al., 2021, 2023) downsample activations in intermediate layers to reduce computation and enable longer contexts. COLT5 (Ainslie et al., 2023) proposes conditional computation to save memory and enable larger contexts. Memorizing Transformer (Wu et al., 2022) uses kNN lookup to pick up the most relevant tokens, which might also be seen as a way to reduce the computational complexity of attention. Our work adheres to this approach and aims to train a key space that handles longer attention context length (e.g., by mitigating the distraction issue) and, thus, has better long-context capabilities.
Fine-tuning LLMs for longer retrieval
@@
Prior works such as RETRO (Borgeaud et al., 2022) (RETROfitting) and Memorizing Transformer (Wu et al., 2022) have demonstrated a promising path for fine-tuning existing LMs to add new capabilities without the need to retrain the entire model. In contrast to those approaches our method is not framed as a retrieval but as a way of extending the context of the model. In contrast to RETRO, we propose a single-stage method for context extension instead of a two-stage retrieve-then-embed approach. We provide a more detailed comparison with the Memorizing Transformer in Appendix C.3. More recently, a number of works have explored fine-tuning LLaMA to extend its context length. Landmark attention (Mohtashami and Jaggi, 2023) proposes a compression scheme of LLM’s context into landmarks, increasing the context length of LLaMA-7B to $32K$. Position Interpolation (PI, (Chen et al., 2023) and (kaiokendev, 2023)) introduces a modification to the rotary positional encoding scheme that enables fine-tuning for $32K$ context. In contrast to this work, our method does not rely on positional encodings, following the findings from (Haviv et al., 2022). Removing positional encoding in additional context allows us to extrapolate to $256k$ tokens, although the model was only trained on sequences up to $8K$, yielding theoretically unbounded context length.
Zero-shot methods
KNN-LM (Khandelwal et al., 2019) shows that one can improve the performance of a LLM by combining two probability distributions. One created by a pre-trained model, and one based on the similarity between the embedding of the currently processed token and the embeddings of tokens retrieved from a large database. Meanwhile, we extend the model context in a subset of attention layers, potentially allowing for reasoning within this extended context. Parallel Context Windows for Large Language Models (Ratner et al., 2023) introduces a method for extending the context of language models without training. They achieve this by embedding several context windows independently in parallel and allowing only a subset of tokens to attend to all windows. On the other hand, we fine-tune existing models and allow all tokens to attend to all previous tokens but only in a subset of layers. Additionally, our method allows us to improve the structure of the key-value space of the existing models.
@@
Contrastive learning aims to learn good representations by comparing positive and negative examples. CLIP (Radford et al., 2021) and SimCLR (Chen et al., 2020) are two popular contrastive learning methods that have achieved state-of-the-art performance in the image domain. During contrastive pre-training, negative examples are kept in the same batch to learn to distinguish them from positive examples. Scaling the batch size in contrastive learning has been demonstrated to enhance the quality of representations, as shown in (Gao et al., 2021b). It has been suggested (Gao et al., 2019) that the embedding space in language modeling suffers from degeneracy, where embeddings are tightly packed in a narrow cone, making it difficult to distinguish between them. TRIME (Zhong et al., 2022) proposes a training approach designed for training LMs with memory augmentation, which uses negatives to improve the quality of representations. The main difference between this and our approach is that we incorporate negatives into the chosen subset of attention layers instead of interpolating in the output layer and use the standard language modeling loss. TRIME (Zhong et al., 2022) also focuses on retrieval from large databases, whereas we focus on extending the context of the model. ContraCLM (Jain et al., 2023) applies contrastive losses at both the token and sequence levels during training to promote more uniformly distributed, isotropic representations. It is shown to enhance the discrimination of representations on textual semantic similarity benchmarks. While ContraCLM focuses on improving the general expressiveness of representations, our work introduces contrastive-inspired techniques designed specifically for training the attention mechanism to handle longer context lengths. Nonetheless, exploring other contrastive learning objectives could be beneficial for further improving the key structure in future work.
@@ 3 FoT: Focused Transformer
Our method, the Focused Transformer (FoT), is a simple plug-and-play extension of transformer models and can be used both to train new models or fine-tune existing, possibly large, models with longer context. To this end, FoT uses memory attention layers and the crossbatch training procedure. Memory attention layers enable the model to retrieve information from the additional context at inference time, effectively extending the context. The crossbatch training procedure biases the model to learn $(key,value)$ representations, which are easy to use by a memory attention layer. See Figure 2 for an overview of the FoT architecture and Appendix L for pseudocode.
Figure 2: The Focused Transformer overview. During inference, a memory attention layer (green) uses additional context of $(key,value)$ pairs via kNN lookup, which effectively extends its context length. This layer is trained using crossbatch. Namely, the tokens from the current context $C_{curr}$ attend in a differentiable way (Att + $\nabla$) to the previous context $C_{prev}$ of the same document and, importantly, $d-1$ contexts of other documents. The latter serve as ’negative’ examples intended to better shape the $(key,value)$ space.
@@ 3.1 Memory attention layers
Memory attention layers $\mathcal{L}$ are endowed with access to an additional context during inference. Namely, each query in $\ell\in\mathcal{L}$ attends to preceding keys from the local context and the top $k$ most matching keys (i.e. having the largest inner product with the query) from memory. The memory keys are ranked by the inner product with the query and retrieved using the kNN search algorithm. We use the exact kNN search implemented in FAISS (Johnson et al., 2017). The memory is populated incrementally with $(key,value)$ pairs processed by $\ell$ beforehand. Our memory attention layer design is closely related to (Wu et al., 2022), we follow most of its design choices, except for the gating, which we replace with a simpler mechanism, which turns out to be more effective in our applications. See details in Section C.3 and Appendix B.2. We remove positional encodings in memory layers in all our models except LongLLaMAs. This allows LongLLaMA checkpoints to be a drop-in replacement for LLaMA checkpoints. We treat the kNN search algorithm as an approximation of full dense attention, which opens the doors for future speed-ups.
@@ 3.2 Crossbatch training procedure
The operation is fully differentiable, and thus, we improve all the $(key,value)$ pairs in $p^{\delta}$. Two, the procedure is easy to implement; it does not require any additional loss (i.e., uses the standard transformer training objective) and is done on the level of the data loading pipeline and a minor self-attention change. The only new hyperparameter is $d$, which prescribes the ratio of positive to negative samples. Typically, we find it beneficial to start with small $d\leq 8$ (otherwise, the model tends to ignore the previous local context) and later switch to bigger values, say $d\geq 64$. Appendix B.3 provides more details about the method. Listing 1 outlines an implementation of the crossbatch.
@@ 3.3 The distraction issue
In this section, we conceptualize what we call the distraction issue and hypothesize it is one of the key problems in dealing with long multi-document contexts (like large code repositories). Namely, during the standard training, the model is not incentivized to distinguish the keys from different documents. We measure that the attention mass is evenly spread on the related and unrelated documents; see Figure 3. More precisely, for a document $\delta$, let $w_{ij}$ be the softmax weights related to $p^{\delta}_{ij}$ constructed as described in Section 3.2. We define the positive attention mass as $r_{d}:=\sum_{j}w_{1j}/\sum_{i=1}^{d}\sum_{j}w_{ij}$. We observe that $r_{d}\approx 1/d$, which can be interpreted as the fact that the attention is equally distracted by the positive (coming from the current document at $i=1$) and negative keys. This is an undesirable property since when scaling the memory, the attention becomes increasingly distracted. We show that the crossbatch mostly alleviates the distraction issue, resulting in a focused attention. More information can be found in Appendix B.4. In Section 5.3, we also show that the distraction issue has a harmful effect on metrics like perplexity.
@@ 4.2 Context length extrapolation on the passkey retrieval task
4.3 Question answering over research papers
In Table 6 we present the performance on the validation set of Qasper (Dasigi et al., 2021) from SCROLLS (Shaham et al., 2022) and compare our results to LongChat 7B (Ma and Zhang, 2023) and two baseline short-context models. We note that our model shows gains from increased context length.
@@ 4.6 Performance on short-context tasks
Fine-tuning for longer contexts could hurt performance on the original context length ($2K$), as the training data distribution changes. We show that this is not the case for the LongLLaMA models by evaluating them using the LM Evaluation Harness library (Gao et al., 2021a). On most tasks, the performance is kept intact; see Appendix A.4 for details. This also confirms that LongLLaMAs could be used as a drop-in replacement of LLaMA models as they are compatible with the original LLaMA inference code.
@@ 5.1 Experimental setup
Architecture For experiments described in this section we use decoder-only Transformer (Vaswani et al., 2017) models with $12$ layers and $184M$ parameters (unless stated otherwise). Following Wu et al. (2022); we pick $\ell=8$ as the memory attention layer. We tune $k=128$, the number of top keys retrieved by kNN. In most experiments, we start training with a small crossbatch dimension $d\leq 8$ and switch to $d\geq 64$ after some training. For more details about the architecture and hyperparameters, see Appendix B and Appendix E.
Evaluation We distinguish two evaluation settings: single-document (abbreviated to single-doc) and multi-document (abbreviated to multi-doc). The single-doc setting is typically used for evaluating models that process long contexts. Here, we clear the memory for each new document, ensuring that only the current document is available in the context. The multi-doc setting retains memory across multiple documents without resets. This scenario tests whether the model can ignore irrelevant information and focus on the relevant data, which can be useful in setups like repository-level code generation.
Datasets We evaluate on the following long-context language modeling datasets: PG-19 (English books), arXiv (mathematical papers), GitHub (code), and Isabelle (formal proofs). PG-19 (Rae et al., 2019) is a large dataset of English-language books published prior to 1919, sourced from the Project Gutenberg archive. This dataset is a well-established benchmark for evaluating long-context language models (Sun et al., 2021). The arXiv dataset contains source of papers labeled as "Mathematics" that were obtained by downloading articles through the arXiv Bulk Data Access. The token count per paper in this dataset is comparable to that of a book in PG19. For details on the remaining datasets, refer to Appendix H.